# MDD-5k: A New Diagnostic Conversation Dataset for Mental Disorders Synthesized via Neuro-Symbolic LLM Agents

### v1

@@ Acknowledgments
We are grateful for the GPU resources and accessible OpenAI API key provided by Shanda Group. Chen Frontier Lab for AI and Mental Health, Tianqiao and Chrissy Chen Institute leads the project of collecting mental disorders patient cases by cooperating with Shanghai Mental Health Center. We also appreciate the support and discussion from psychiatrists in Shanghai Mental Health Center.
@@ Appendix A Pseudo-code of the Proposed Framework
In this section, we detailed explain the multi-agent framework for diagnostic conversation simulation through pseudo-code. As illustrated in Algorithm 1, the doctor agent, patient agent and tool agent are initialized with large language models. The whole process is guided by the tool agent. When the diagnostic conversation finishes, the diagnosis result from patient case will be appended to the dialogue history as label.
@@
Appendix B Case Study
We present three cases of the synthesized diagnostic conversations in MDD-5k, as well as all the material for carrying out the conversation simulation, including patient case, doctor prompt, diagnostic topic shift, and generated fictitious patient experience. Diagnostic conversation 1 and 2 are generated with the same patient case 1 to highlight our one-to-many patientcase-to-dialogue framework is capable of generating diverse conversation. Dialogue conversation 3 is generate with another patient case 2. The red words are diagnosis result, namely the label for each conversation. Both Chinese and English samples are provided. As to the Chinese version, Figure 4, Figure 8 and Figure 11 show the patient cases and other related material. Figure 5 to Figure 7 show diagnostic conversation 1. Figure 9 and Figure 10 show diagnostic conversation 2. Figure 12 to Figure 14 show diagnostic conversation 3. As to the English version, Figure 15 and Figure 16, Figure 22 and Figure 23, Figure 28 and Figure 29 show the patient cases and other related material. Figure 17 to Figure 21 show diagnostic conversation 1. Figure 24 to Figure 27 show diagnostic conversation 2. Figure 30 to Figure 34 show diagnostic conversation 3.
Appendix C Feedback from Psychiatrists
During human evaluation, psychiatrists also provide us feedback for the synthesized diagnostic conversation. The discussion between professional psychiatrists and us is listed.
  • The patient response is so “sound and perfect”. In real diagnostic conversation, many mental disorders patients struggle to respond to the psychiatrists questions. For this issue, we are not sure whether a perfect patient answer or a poor but similar to real situation answer will benefit from the perspective of natural language processing.
  • The proportion of patients suffer from major depressive disorder is too small. The number of pediatric patients is also small.
  • Lack the role of companion (e.g. parent) which is quite common in real scenarios. This actually inspires us to design a multi-party dialogue system for dealing with diagnostic conversation simulation in the future.
@@ Appendix C Feedback from Psychiatrists
Figure 4: Chinese version of patient information and doctor persona for diagnostic conversation 1.
Figure 5: Part 1 of the Chinese synthesized diagnostic conversation 1.
Figure 6: Part 2 of the Chinese synthesized diagnostic conversation 1.
Figure 7: Part 3 of the Chinese synthesized diagnostic conversation 1.
Figure 8: Chinese version of patient information and doctor persona for diagnostic conversation 2.
Figure 9: Part 1 of the Chinese synthesized diagnostic conversation 2.
Figure 10: Part 2 of the Chinese synthesized diagnostic conversation 2.
Figure 11: Chinese version of patient information and doctor persona for diagnostic conversation 3.
Figure 12: Part 1 of the Chinese synthesized diagnostic conversation 3.
Figure 13: Part 2 of the Chinese synthesized diagnostic conversation 3.
Figure 14: Part 3 of the Chinese synthesized diagnostic conversation 3.
Figure 15: English version of patient information and doctor persona for diagnostic conversation 1.
Figure 16: English version of patient information and doctor persona for diagnostic conversation 1.
Figure 17: Part 1 of the English synthesized diagnostic conversation 1.
Figure 18: Part 2 of the English synthesized diagnostic conversation 1.
Figure 19: Part 3 of the English synthesized diagnostic conversation 1.
Figure 20: Part 4 of the English synthesized diagnostic conversation 1.
Figure 21: Part 5 of the English synthesized diagnostic conversation 1.
Figure 22: English version of patient information and doctor persona for diagnostic conversation 2.
Figure 23: English version of patient information and doctor persona for diagnostic conversation 2.
Figure 24: Part 1 of the English synthesized diagnostic conversation 2.
Figure 25: Part 2 of the English synthesized diagnostic conversation 2.
Figure 26: Part 3 of the English synthesized diagnostic conversation 2.
Figure 27: Part 4 of the English synthesized diagnostic conversation 2.
Figure 28: English version of patient information and doctor persona for diagnostic conversation 3.
Figure 29: English version of patient information and doctor persona for diagnostic conversation 3.
Figure 30: Part 1 of the English synthesized diagnostic conversation 3.
Figure 31: Part 2 of the English synthesized diagnostic conversation 3.
Figure 32: Part 3 of the English synthesized diagnostic conversation 3.
Figure 33: Part 4 of the English synthesized diagnostic conversation 3.
Figure 34: Part 5 of the English synthesized diagnostic conversation 3.

### v2

@@ Acknowledgments
We are grateful for the GPU resources and accessible LLM API keys provided by Shanda Group. Chen Frontier Lab for AI and Mental Health, Tianqiao and Chrissy Chen Institute leads the project of collecting mental disorders patient cases by cooperating with Shanghai Mental Health Center. We also appreciate the support and discussion from psychiatrists in Shanghai Mental Health Center.
@@ Appendix A Pseudo-code of the Proposed Framework
We detailedly explain the multi-agent framework for diagnostic conversation simulation through pseudo-code. As illustrated in Algorithm 1, the doctor agent, patient agent and tool agent are initialized with large language models. The whole process is guided by the tool agent. When the diagnostic conversation finishes, the diagnosis result and treatment opinion from patient case will be appended to the dialogue history as label.
@@
Appendix B Feedback from Psychiatrists
During human evaluation, psychiatrists also provide us feedback for the synthesized diagnostic conversation. The discussion between professional psychiatrists and the authors is listed.
  • The patient response is too “sound and perfect”. In real diagnostic conversation, many mental disorders patients struggle to respond to the psychiatrists questions. For this issue, we are not sure whether a perfect patient answer or a poor but similar to real situation answer will benefit from the perspective of natural language processing.
  • The proportion of patients suffer from severe depressive disorder is too small. The number of pediatric patients is also small.
  • The synthesized dialogue lacks the presence of a companion role (e.g. parent), which is quite common in real scenarios. This actually inspires us to design a multi-party dialogue system for dealing with diagnostic conversation simulation in the future.